I am in intermediate algebra college level.. The author of the book is Tussy & Gustafson..
Slope: Equations of Lines: Systems of Equations
1) Find the slope of the line that passes through the points (2,1) and (0,4).
2) Find the equation, in slope-intercept form, of the line that passes through the points (-3,-7)and (-2,8).

Answer by Nate(3500)

1) Where the points: (x1,y1) and (x2,y2)
Slope = (y2 - y1)/(x2 - x1) = (4 - 1)/(0 - 2) = -3/2
2)
Slope = (y2 - y1)/(x2 - x1) = (8 + 7)/(-2 + 3) = 15/1
point-slope form:
y - y1 = m(x - x1)
y + 7 = 15(x + 3)
y + 7 = 15x + 45
y = 15x + 38